ADVANCED ANALYSIS Suppose that the linear equation for consumption in a hypothetical economy is: C= 60 + 0.7Y Also suppose thc. Consumption d. APC = e. Saving = 1. APS =
Answer #1

c = 60+0.74 Y = $600. mpe = dc dy =) MPC=0.7 So, IMPc = 0.7 b) mps I- MPC = 1-0-7= 0.3 & Imps = 0.3 c) consumption : 60 to 7